A basketball is shot at 12.0 m/s at a 55.0 degree angle. What is the velocity of the ball 2.16 seconds later?
1
Expert's answer
2020-05-14T09:12:51-0400

The ball has two projections of speed: "v_{x}" and "v_{y}" .

At the initial moment:

"v_{0x}=v_{0}\times \cos(\alpha) \newline
v_{0y}=v_{0}\times \sin(\alpha)"

The x-projection of speed will not change in time.

The y-projection:

"v_{y}=v_{0y}+at=v_{0y}-gt"

Take "g=10" :

"v_{x}=v_{0x}=12\times\cos(55)\approx6.88\newline
v_{y}=12\times\sin(55)-10\times2.16\approx-11.76"

So:

"v=\sqrt{v_{x}^2+v_{y}^2}=\sqrt{47.3344+138.2976}\approx13.63m/s"